OpenStudy (anonymous):

Convert the polar equation to a rectangular equation. Then determine the graph's slope and y-intercept: rsin(θ-π/4)=2

OpenStudy (anonymous):

Two things you have to know for this problem: 1. x= rcosθ y= rsinθ ultimately you want to get the equation in this shape!! 2. Indentity: sin(a+b)= sin(a)cos(b)+cos(a)sin(b)--- Use this to open up the sin(θ-π/4). Because we have a minus θ-π/4 here we will have a minus sin(a)cos(b)-cos(a)sin(b) here.

OpenStudy (anonymous):

so the answer is y=-x+2root2; slope=1; y-intercept is 2root2 but my chemistry teacher (who also teaches calculus and other math courses) got 2 and im confused.he did it they way you said

OpenStudy (anonymous):

yeah, i got 2root2 too, hahah. the slope is 1, not -x! He probably made a mistake somewhere, or we did

OpenStudy (anonymous):

nope im sorry i was wrong my friend just texted me and it turns out my teacher used pi/2 instead of pi/4!

OpenStudy (anonymous):

hahaha, does help to have the right question :)